$SPAD/src/input exlap.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool exlap.output
)set message test on
)set message auto off
)clear all
--S 1 of 6
laplace(2/t * (1 - cos(a*t)), t, s)
--R
--R
--R             2    2
--R   (1)  log(s  + a ) - 2log(s)
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 6
laplace((exp(a*t) - exp(b*t))/t, t, s)
--R
--R
--R   (2)  - log(s - a) + log(s - b)
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 6
laplace(exp(a*t+b)*Ei(c*t), t, s)
--R
--R
--R          b    s + c - a
--R        %e log(---------)
--R                   c
--R   (3)  -----------------
--R              s - a
--R                                                    Type: Expression(Integer)
--E 3
)clear all
--S 4 of 6
laplace(a*Ci(b*t) + c*Si(d*t), t, s)
--R

--R
--R               2    2
--R              s  + b             d
--R        a log(-------) + 2c atan(-)
--R                  2              s
--R                 b
--R   (1)  ---------------------------
--R                     2s
--R                                                    Type: Expression(Integer)
--E 4
)clear all
--S 5 of 6
laplace(sin(a*t) * cosh(a*t) - cos(a*t) * sinh(a*t), t, s)
--R
--R
--R             3
--R           4a
--R   (1)  --------
--R         4     4
--R        s  + 4a
--R                                                    Type: Expression(Integer)
--E 5
)clear all
--S 6 of 6
laplace(t**4 * exp(-a*t) / factorial(4), t, s)
--R
--R
--R                            1
--R   (1)  ----------------------------------------
--R         5       4      2 3      3 2     4     5
--R        s  + 5a s  + 10a s  + 10a s  + 5a s + a
--R                                                    Type: Expression(Integer)
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5